# InstructRAG: Instructing Retrieval-Augmented Generation via Self-Synthesized Rationales

### v2

@@ 3.4 Analysis
InstructRAG-ICL and InstructRAG-FT generalize well to unseen tasks. Figure 4 demonstrates the generalization ability of our method in both training-free and trainable settings. For the in-domain (ID) method, it directly utilizes target domain demonstrations (in training-free settings) or is trained on the target domain task (in trainable settings). In contrast, the out-of-domain (OOD) method can only learn from demonstrations or training data in the source domain, and have no prior knowledge of the target domain. In this case, the model must leverage the knowledge learned from the source domain task to solve the unseen target domain task. The results show that our method (i.e., InstructRAG-ICL and InstructRAG-FT) consistently outperforms the baselines across various scenarios in both in-domain and out-of-domain settings, demonstrating strong task generalizability. One counter-intuitive finding is that in the scenario of generalizing from long-form to short-form QA task (Figure 4(b)), the training-free OOD method substantially outperforms its in-domain counterpart. We speculate that the training-free OOD method achieves better performance because it benefits from the demonstrations with long answers from the source domain (ASQA). The reason is that the questions in ASQA are ambiguous and can have multiple interpretations, and ground-truth long answers often address the questions from various perspectives, which can be regarded as a form of chain-of-thought demonstration.
@@
Furthermore, we also study the generalizability of InstructRAG to a non-QA knowledge-intensive task such as code generation. As presented in Table 5(a), we directly apply InstructRAG-FT trained on the QA task (PopQA) to solve the unseen code generation task (HumanEval [12]), following the CodeRAG-Bench setup [87]. We evaluate the code generation performance using the standard pass@k metric and compare our method with the off-the-shelf Llama-3-8B-Instruct as the baseline. It can be observed that our method consistently achieves better generalization performance in the unseen code generation task in both non-retrieval and RAG settings. This finding aligns with our observation that InstructRAG trained on QA tasks tends to generate more text-based comments that articulate the design of coding solutions compared to the off-the-shelf Llama-3-8B-Instruct, thereby leading to more accurate code generation.
Evaluation with LLM-as-a-judge. Despite being standard evaluation metrics for question-answering, accuracy or exact match are known to be imperfect [13] as they mainly rely on pattern-matching to judge whether the predicted answer aligns with the ground-truth answer. Such metrics cannot handle cases where the predicted answer and ground-truth answer are synonyms (e.g., “Donald Trump” vs “Donald J. Trump” cannot be correctly recognized as a match), leading to biased evaluation results.
Therefore, we adopt the LLM-as-a-judge [7, 115] approach to evaluate whether the model response aligns with the ground-truth answer using GPT-4o [66], which allows the judge to consider semantic equivalence and is expected to yield a more fair evaluation. As shown in Table 5(b), we evaluate our method and baseline models on the open-domain Natural Questions benchmark in both training-free and trainable RAG settings. Compared to pattern-matching based metrics, LLM-as-a-judge generally leads to higher evaluation results, mostly due to its capability to accurately match semantically equivalent phrasings. Notably, our method consistently outperforms baselines under both pattern-matching based and LLM-based evaluation metrics, further validating the effectiveness of InstructRAG.
@@ 4.1 Retrieval-augmented Generation
Retrieval-augmented generation (RAG) is a widely adopted approach to enhance large language models (LLMs) with external knowledge [2, 4, 5, 19, 23, 42, 52, 70, 74], demonstrating promising potential to reduce hallucinations and enhance the generation accuracy of LLMs across various real-world applications [9, 37, 55, 58, 77, 80, 117]. Recently, a growing research effort has been devoted to enhancing RAG from various aspects, such as improving decoding efficiency [64, 36, 57], exploring long-context retrieval [96, 104], compressing prompts [32, 95], and addressing practical concerns such as adversarial retrieval [92, 116, 118] and privacy leakage [26, 108]. Despite their advantages, these RAG systems inevitably suffer from irrelevant information introduced by imperfect retrievers or noisy retrieval corpora. However, most existing works typically address this issue by improving the retrieval quality and reducing noise exposure to the model [22, 33, 72, 86, 100, 102, 109, 111]. Notable methods include adaptive retrieval [3, 35, 103] and query rewriting [8, 62]. In contrast, our work focuses on an orthogonal direction of developing explicit denoising methods for RAG, thereby enhancing the model’s noise robustness and generation accuracy, even in highly noisy contexts.
@@ 4.2 Eliciting Reasoning in Large Language Models
Recent studies have extensively explored the reasoning capability of LMs, but typically not in the context of RAG where potentially noisy retrieved contents may mislead the reasoning if not properly addressed. Chain-of-thought (CoT) prompting [89] is an effective method to elicit step-by-step reasoning from LMs by showing exemplars with detailed explanations (i.e., rationales [16, 49, 69, 107]) that lead to the final answer. However, such works often requires manually crafted demonstrations [84, 97], which is costly and requires extensive efforts and domain knowledge [114]. To mitigate this limitation, automatic chain-of-thought prompting (Auto-CoT) [112] is introduced to automatically select instances from the corpus coupled with zero-shot CoT [46], where the rationales are generated by LMs. Furthermore, it has been shown that CoT reasoning can be elicited even without explicit prompting, particularly for instruction-tuned LMs [85]. Another related work shows rationales generated by small models can help large models reason better [50]. Although rationalization has been extensively investigated in many NLP tasks [10, 20, 68, 90], none of them are designed for RAG, and how to leverage the instruction-following abilities of LMs for explicit denoising in the context of RAG is still underexplored.
@@ 5 Conclusion
In this work, we presented InstructRAG, a simple retrieval-augmented generation (RAG) approach that explicitly denoises retrieved contents and produces accurate generations. By leveraging the strong instruction-following abilities of large language models, InstructRAG generates detailed rationales that articulate how the ground-truth answers can be derived from the retrieved documents. These synthetic rationales can serve as either in-context learning examples or supervised fine-tuning data, enabling the model to learn an explicit denoising process. Experiments on five knowledge-intensive benchmarks show InstructRAG consistently outperforms state-of-the-art RAG approaches with significant improvements in both training-free and trainable settings. Compared to the best baseline method, InstructRAG achieves an average improvement of 8.3% across all benchmarks, demonstrating its effectiveness in enhancing the noise robustness of retrieval-augmented generation.
@@
In this work, we mainly conduct experiments on question answering-type tasks, and it remains unclear how our method may generalize to other scenarios (e.g., open-ended generation). Moreover, despite being the standard evaluation metrics, both accuracy and exact match are biased and cannot perfectly reflect the quality of the model’s generations. For instance, such metrics heavily rely on string matching, which assesses correctness at the lexical level rather than the semantic level, thereby failing to recognize different phrasings that convey identical meanings. The evaluation results also suffer from length bias, as longer generations tend to achieve higher accuracy. Exploring more advanced metrics like using LLMs as judges would better evaluate RAG model generations [102]. Another potential limitation is that our model might be subject to sample bias in the training data. Incorporating bias mitigation methods [17, 41, 101] would be helpful for further improving our work.
@@
Future research directions include exploring more advanced techniques for generating high-quality rationales, such as incorporating domain-specific knowledge or leveraging multi-task learning to enable better generalization across various tasks. For instance, although the consistency ratio between synthetic rationales and ground-truth answers on training samples with at least one relevant document achieves 98%, the overall consistency ratio on all training samples is only 89%. This is because for some samples, none of the retrieved documents is relevant to the question, which significantly compromises the quality of the generated rationales. Therefore, it will be interesting to fully explore the potential of our method by incorporating additional designs such as a filtering mechanism, which we leave as future work. Additionally, investigating the scalability of our method to larger datasets and models could further demonstrate its generalizability and robustness. It will also be interesting to evaluate the model performance under long-context settings with a dynamic or extremely large number of retrieved documents. Finally, integrating our method with other advanced retrieval techniques, such as active retrieval, could potentially lead to even better performance on knowledge-intensive tasks.
Acknowledgments
The authors would like to thank Xinyu Zhu from University of Virginia, Tianyu Gao and Zexuan Zhong from Princeton NLP group for their valuable feedback and discussions.
Appendix A Implementation Details
@@
Retrieval setup. Following [3, 70], we use the Wikipedia dump from [39] as the external retrieval corpus for all five benchmarks studied in this work, where each document is a disjoint text block of up to 100 words extracted from a Wikipedia article. We compared all RAG methods under a diverse retrieval environment with various sparse and dense retrievers and number of retrieved documents. Specifically, we use Contriever-MS MARCO as the retriever for PopQA and TriviaQA, DPR for Natural Questions, GTR for ASQA, and BM25 for 2WikiMultiHopQA. By default, we retrieve the top 5 documents from the retrieval corpus for each query in all tasks except 2WikiMultiHopQA, where the top 10 documents are retrieved. We use the official weights for all dense retrievers and the implementation from Pyserini [54] for the sparse retriever BM25.
@@
Training details. Our models are trained on 4 Nvidia H100 GPUs with 80GB memory via full-parameter fine-tuning. We use fully sharded data parallelism (FSDP) for distributed training, along with FlashAttention [14] and bf16 mixed precision training enabled for computation efficiency. By default, all models are trained using the Adam optimizer [45] for 2 epochs, with a batch size of 128, a learning rate of 2.5e-5, and a cosine learning rate schedule with 3% warmup steps. For the trainable baseline vanilla SFT, we use a slightly different learning rate of 2e-5 based on our hyper-parameter search results. To fairly compare with Self-RAG and RetRobust, we re-implement them using Llama-3-Instruct-8B. We also optimize their performance through an extensive hyper-parameter search with learning rates in [8e-6, 1e-5, 2e-5] and training epochs in [1, 2, 3]. For Self-RAG, we use a learning rate of 1e-5 with a single training epoch. For RetRobust, we use a learning rate of 2e-5 with two training epochs. The only exception is the training for RetRobust on 2WikiMultiHopQA, where we train the model for 5 epochs on the augmented training set released by the original authors. The maximum token length for all models is fixed at 4096.
@@
Inference details. By default, the number of demonstrations used in InstructRAG-ICL and the baseline method few-shot demonstration with instruction is set to be 2. We use vLLM [48] to load models for memory-efficient inference and adopt the greedy decoding strategy for model generation.
Appendix B Case Study
@@
Attention visualization. To intuitively understand the denoising process of our InstructRAG, we visualize its attention from the answer to retrieved documents. As pointed out by a recent work [106], only attention distributions from deep layers can accurately reflect the LM’s retrieval behavior and focus on key information, while attention from shallow layers usually do not imply meaningful patterns. Therefore, we only plot the attention weights of the last 10 layers (Layer 22 to Layer 31). As presented in Figure 5, our model accurately identifies the only benign document from noisy input, showing a strong denoising signal compared to vanilla SFT.
@@
Appendix C Prompt Templates
@@
Rationale generation. Below are the prompt templates for rationale generation used in all five tasks. Table 6 shows the rationale template used in the ablation study. For simplicity, we use the same prompt structure (Table 7) for all tasks with minor differences in task-specific instructions (Table 8).

### v3

@@ 3.4 Analysis
InstructRAG-ICL and InstructRAG-FT generalize well to unseen tasks. Figure 4 demonstrates the generalization ability of our method in both training-free and trainable settings. For the in-domain (ID) method, it directly utilizes target domain demonstrations (in training-free settings) or is trained on the target domain task (in trainable settings). In contrast, the out-of-domain (OOD) method can only learn from demonstrations or training data in the source domain, and have no prior knowledge of the target domain. In this case, the model must leverage the knowledge learned from the source domain task to solve the unseen target domain task. The results show that our method consistently outperforms the baselines across various scenarios in both in-domain and out-of-domain settings, demonstrating strong task generalizability. One counter-intuitive finding is that in the scenario of generalizing from long-form to short-form QA task (Figure 4(b)), the training-free OOD method substantially outperforms its in-domain counterpart. We speculate that the training-free OOD method achieves better performance because it benefits from the demonstrations with long answers from the source domain (ASQA). The reason is that the questions in ASQA are ambiguous and can have multiple interpretations, and ground-truth long answers often address the questions from various perspectives, which can be regarded as a form of chain-of-thought demonstration.
@@
Furthermore, we also study the generalizability of InstructRAG to a non-QA knowledge-intensive task such as code generation. As presented in Table 5(a), we directly apply InstructRAG-FT trained on the QA task (PopQA) to solve the unseen code generation task (HumanEval (Chen et al., 2021)), following the CodeRAG-Bench setup (Wang et al., 2024c). We evaluate the code generation performance using the standard pass@k metric and compare our method with the off-the-shelf Llama-3-8B-Instruct as the baseline. It can be observed that our method consistently achieves better generalization performance in the unseen code generation task in both non-retrieval and RAG settings. This finding aligns with our observation that InstructRAG trained on QA tasks tends to generate more text-based comments that articulate the design of coding solutions compared to the off-the-shelf Llama-3-8B-Instruct, thereby leading to more accurate code generation.
@@
Evaluation with LLM-as-a-judge. Despite being standard evaluation metrics for question-answering, accuracy or exact match are known to be imperfect (Cuconasu et al., 2024) as they mainly rely on pattern-matching to judge the correctness of model predictions. Such metrics cannot handle cases where the prediction and ground-truth are synonyms (e.g., “Donald Trump” vs “Donald J. Trump” cannot be correctly recognized as a match), leading to biased evaluation results. Therefore, we use LLM-as-a-judge (Bubeck et al., 2023; Zheng et al., 2024b) to evaluate the predictions with GPT-4o (OpenAI, 2024), which allows the judge to consider semantic equivalence and is expected to yield a more fair evaluation. As shown in Table 5(b), we evaluate our method and baseline models on the open-domain Natural Questions benchmark in both training-free and trainable RAG settings. Compared to pattern-matching based metrics, LLM-as-a-judge generally leads to higher evaluation results, mostly due to its capability to accurately match semantically equivalent phrasings. Notably, our method consistently outperforms baselines under both pattern-matching based and LLM-based evaluation metrics, further validating the effectiveness of InstructRAG.
@@ 4.1 Retrieval-augmented Generation
Retrieval-augmented generation (RAG) is a widely adopted approach to enhance large language models (LLMs) with external knowledge (Asai et al., 2023a; 2024; Borgeaud et al., 2022; Gao et al., 2023b; Guu et al., 2020; Khandelwal et al., 2019; Lewis et al., 2020; Ram et al., 2023; Shi et al., 2023), demonstrating promising potential to reduce hallucinations and enhance the generation accuracy of LLMs across various real-world applications (Chase, 2022; Jin et al., 2024b; Liu, 2022; Lu et al., 2022; Siriwardhana et al., 2023; Tan et al., 2024; Zhou et al., 2022; Liu et al., 2024b; Xiong et al., 2025). Recently, a growing research effort has been devoted to enhancing RAG from various aspects, such as improving decoding efficiency (Merth et al., 2024; Jin et al., 2024a; Liu et al., 2023; Wang et al., 2024b), exploring long-context retrieval (Xu et al., 2024a; Yen et al., 2024), compressing prompts (Jiang et al., 2023a; Xu et al., 2023; Cheng et al., 2024), and addressing practical concerns such as adversarial retrieval (Xiang et al., 2024; Zhong et al., 2023; Zou et al., 2024) and privacy leakage (Huang et al., 2023b; Zeng et al., 2024). Despite their advantages, these RAG systems inevitably suffer from irrelevant information introduced by imperfect retrievers or noisy retrieval corpora. However, most existing works typically address this issue by improving the retrieval quality and reducing noise exposure to the model (Gupta et al., 2024; Jiang et al., 2024a; Sarthi et al., 2024; Wang et al., 2024a; Yan et al., 2024; Yang et al., 2024; Zhang et al., 2024a; b). Notable methods include adaptive retrieval (Asai et al., 2023b; Jiang et al., 2023b; Yao et al., 2022) and query rewriting (Chan et al., 2024; Mao et al., 2024). In contrast, our work focuses on an orthogonal direction of developing explicit denoising methods for RAG, thereby enhancing the model’s noise robustness and generation accuracy, even in highly noisy contexts.
@@ 4.2 Eliciting Reasoning in Large Language Models
Recent studies have extensively explored the reasoning capability of LMs, but typically not in the context of RAG where potentially noisy retrieved contents may mislead the reasoning if not properly addressed. Chain-of-thought (CoT) prompting (Wei et al., 2022) is an effective method to elicit step-by-step reasoning from LMs by showing exemplars with detailed explanations (i.e., rationales (Feng et al., 2024; Lampinen et al., 2022; Rajani et al., 2019; Zelikman et al., 2024)) that lead to the final answer. However, such works often requires manually crafted demonstrations (Wang et al., 2022; Xu et al., 2024b), which is costly and requires extensive efforts and domain knowledge (Zheng et al., 2024a). To mitigate this limitation, other methods have been introduced to automatically select instances from the corpus (Zhang et al., 2022) or curate demonstrations by the LM itself (Chen et al., 2023a), coupled with zero-shot CoT (Kojima et al., 2022) to generate rationales. Furthermore, it has been shown that CoT reasoning can be elicited even without explicit prompting, particularly for instruction-tuned LMs (Wang & Zhou, 2024). Another related work shows rationales generated by small models can help large models reason better (Lee et al., 2024). Although rationalization has been extensively investigated in many NLP tasks (Chen et al., 2023b; 2022; Ghoshal et al., 2022; Paranjape et al., 2020; Wiegreffe et al., 2021), none of them are designed for RAG, and how to leverage the instruction-following abilities of LMs for explicit denoising in the context of RAG still remains underexplored.
@@ 5 Conclusion
In this work, we presented InstructRAG, a simple retrieval-augmented generation (RAG) approach that explicitly denoises retrieved contents and produces accurate generations. By leveraging the strong instruction-following abilities of large language models, InstructRAG generates detailed rationales that articulate how the ground-truth answers can be derived from the retrieved documents. These synthetic rationales can serve as either in-context learning examples or supervised fine-tuning data, enabling the model to learn an explicit denoising process. Experiments on five knowledge-intensive benchmarks show InstructRAG consistently outperforms state-of-the-art RAG approaches with significant improvements in both training-free and trainable settings. Compared to the best baseline method, InstructRAG achieves an average improvement of 8.3% across all benchmarks, demonstrating its effectiveness in enhancing the noise robustness of retrieval-augmented generation. Limitations and future work are discussed in Appendix A.
Acknowledgments
The authors would like to thank Xinyu Zhu from University of Virginia, Tianyu Gao and Zexuan Zhong from Princeton NLP group for their valuable feedback and discussions. This research was supported in part by the NVIDIA Academic Grant and the OpenAI Researcher Access Program. We thank anonymous reviewers for their constructive and insightful comments.
Reproducibility Statement
To ensure the highest level of reproducibility for our reported results, we have provided:
Complete source code, accessible via the following link: https://github.com/weizhepei/InstructRAG;
Comprehensive implementation details in Appendix B;
All prompt templates used in our experiments in Appendix D.
Appendix A Limitations and Future Works
@@
In this work, we mainly conduct experiments on question answering-type tasks, and it remains unclear how our method may generalize to other scenarios (e.g., open-ended generation). Moreover, despite being the standard evaluation metrics, both accuracy and exact match are biased and cannot perfectly reflect the quality of the model’s generations. For instance, such metrics heavily rely on string matching, which assesses correctness at the lexical level rather than the semantic level, thereby failing to recognize different phrasings that convey identical meanings. The evaluation results also suffer from length bias, as longer generations tend to achieve higher accuracy. Exploring more advanced metrics like using LLMs as judges would better evaluate RAG model generations (Yang et al., 2024). Another potential limitation is that our model might be subject to sample bias in the training data. Incorporating bias mitigation methods (Gallegos et al., 2024; Kawamae, 2023; Yang et al., 2023) would be helpful for further improving our work.
@@
Future research directions include exploring more advanced techniques for generating high-quality rationales, such as incorporating domain-specific knowledge or leveraging multi-task learning to enable better generalization across various tasks. For instance, although the consistency ratio between synthetic rationales and ground-truth answers on training samples with at least one relevant document achieves 98%, the overall consistency ratio on all training samples is only 89%. This is because for some samples, none of the retrieved documents is relevant to the question, which significantly compromises the quality of the generated rationales. Therefore, it will be interesting to fully explore the potential of our method by incorporating additional designs such as a filtering mechanism, which we leave as future work. It will also be interesting to evaluate the model performance under long-context settings with a dynamic or extremely large number of retrieved documents. Finally, integrating our method with other advanced retrieval techniques (Su et al., 2024), such as active retrieval, could potentially lead to even better performance on knowledge-intensive tasks.
Appendix B Implementation Details
@@
Retrieval setup. Following (Asai et al., 2023b; Ram et al., 2023), we use the Wikipedia dump from (Karpukhin et al., 2020) as the external retrieval corpus for all five benchmarks studied in this work, where each document is a disjoint text block of up to 100 words extracted from a Wikipedia article. We compared all RAG methods under a diverse retrieval environment with various sparse and dense retrievers and number of retrieved documents. Specifically, we use Contriever-MS MARCO as the retriever for PopQA and TriviaQA, DPR for Natural Questions, GTR for ASQA, and BM25 for 2WikiMultiHopQA. By default, we retrieve the top 5 documents from the retrieval corpus for each query in all tasks except 2WikiMultiHopQA, where the top 10 documents are retrieved. We use the official weights for all dense retrievers and the implementation from Pyserini (Lin et al., 2021) for the sparse retriever BM25.
@@
Training details. Our models are trained on 4 Nvidia H100 GPUs with 80GB memory via full-parameter fine-tuning. We use fully sharded data parallelism (FSDP) for distributed training, along with FlashAttention (Dao, 2023) and bf16 mixed precision training enabled for computation efficiency. By default, all models are trained using the Adam optimizer (Kingma & Ba, 2014) for 2 epochs, with a batch size of 128, a learning rate of 2.5e-5, and a cosine learning rate schedule with 3% warmup steps. For the trainable baseline vanilla SFT, we use a slightly different learning rate of 2e-5 based on our hyper-parameter search results. To fairly compare with Self-RAG and RetRobust, we re-implement them using Llama-3-Instruct-8B. We also optimize their performance through an extensive hyper-parameter search with learning rates in [8e-6, 1e-5, 2e-5] and training epochs in [1, 2, 3]. For Self-RAG, we use a learning rate of 1e-5 with a single training epoch. For RetRobust, we use a learning rate of 2e-5 with two training epochs. The only exception is the training for RetRobust on 2WikiMultiHopQA, where we train the model for 5 epochs on the augmented training set released by the original authors. The maximum token length for all models is fixed at 4096.
@@
Inference details. By default, the number of demonstrations used in InstructRAG-ICL and the baseline method few-shot demonstration with instruction is set to be 2. We use vLLM (Kwon et al., 2023) to load models for memory-efficient inference and adopt the greedy decoding strategy for model generation.
Appendix C Case Study
@@
Attention visualization. To intuitively understand the denoising process of our InstructRAG, we visualize its attention from the answer to retrieved documents. As pointed out by a recent work (Yu et al., 2023), only attention distributions from deep layers can accurately reflect the LM’s retrieval behavior and focus on key information, while attention from shallow layers usually do not imply meaningful patterns. Therefore, we only plot the attention weights of the last 10 layers (Layer 22 to Layer 31). As presented in Figure 5, our model accurately identifies the only benign document from noisy input, showing a strong denoising signal compared to vanilla SFT.
@@
Appendix D Prompt Templates
@@
Rationale generation. Below are the prompt templates for rationale generation used for all five benchmarks in our work. Table 6 shows the rationale template used in the ablation study (§ 3.3). For simplicity, we use the same prompt structure (Table 7) for all tasks with minor differences in task-specific instructions (Table 8).
@@
Appendix E Example of LLM-as-a-judge
To measure the quality of model-generated rationales, we employ the LLM-as-a-judge approach for a more comprehensive evaluation of the model’s outputs. As illustrated in Table 12, if the rationale is inaccurate despite the final answer being correct (probably due to the use of the LLM’s parametric knowledge), the LLM judge will detect this inconsistency.